Hello,
If i'm travelling at 145 mph,how many miles will I do in 35 minutes?
Many Thanks :)

Answer by malglu(63)

let x be the answer you seek
set this up as two fractions
145/60 = x/35
this is a proportion, so 145 * 35 = 60 * x
145* 35 = 5075.
5075 divided by 60 = 84.5833333
which is the value of x